Waves in a lake are 8.4 m in length and pass an anchored boat once every 2.2 s. What is the speed of the waves?
1
Expert's answer
2020-04-02T10:50:40-0400

The speed of the waves:


"v=\frac{\lambda}{T}"

"v=\frac{8.4}{2.2}=3.8\frac{m}{s}"